Calculate the number of moles present in the following using the formula . number of moles = molar mass 1)40g of Mgo


2)25 of AgNo3






1
Expert's answer
2021-10-26T05:35:40-0400

The molar masses are:

MgO - 40.30 g/mol;

AgNO3 - 169.87 g/mol.


a)

"Moles\ MgO=\frac{mass}{molar\ mass}=\frac{40\ g}{40.30\ g/mol}=0.99\ mol"


b)

"Moles\ AgNO_3=\frac{mass}{molar\ mass}=\frac{25\ g}{169.87\ g/mol}=0.15\ mol"